OpenStudy (anonymous):

Test for convergence or divergence. 1/(n^(1+1/n)) I have no idea where to start with this one..

OpenStudy (anonymous):

Is it \[ \frac {1} {n^{1+\frac 1 n}} \]

OpenStudy (anonymous):

yes thats correct

OpenStudy (anonymous):

I am not sure which method to use here. I was thinking comparison test, but what would i campare it to?

OpenStudy (psymon):

Whoops, took me a sec for the light bulb to click. You have to use the fact that when you multiply exponents that have a common base, the exponents simply get added. The denominator can be rewritten as n^(1)*n^(1/n)

OpenStudy (anonymous):

so could i use the comparison test and use (1/n) and compare it to the whole thing?

OpenStudy (psymon):

That might work, yes. Either way, I would try comparison series as well at first glance.

OpenStudy (psymon):

Well, wait, just choose 1/n^2, since its a convergent p-series.

OpenStudy (anonymous):

so i would compare 1/n^2 to the whole thing?

OpenStudy (psymon):

You can, definitely.

OpenStudy (anonymous):

Use the limit comparison test to see that it is divergent

OpenStudy (anonymous):

\[ \lim_{n->\infty} \frac {\frac{1}{n \ n^{1/n}}} {\frac 1 n}=\frac {1}{n^{1/n}}=1 \]

OpenStudy (anonymous):

Your series behave like \[ \frac 1 n \] which diverges.

OpenStudy (anonymous):

that makes more sense now...thanks Eliass

OpenStudy (anonymous):

YW

OpenStudy (psymon):

Then may I ask what is wrong with using the direct comparison test and using 1/n^2 as a comparison? Just to make sure of why I am wrong.

OpenStudy (anonymous):

TO prove that is divergent you have to say that the n-th term is bigger than a divergent series. 1/n^2 is convergent.

OpenStudy (anonymous):

I have to go to go to an appointment. Enjoy

OpenStudy (psymon):

Alrighty, have fun.

OpenStudy (anonymous):

thanks for your input too Psymon...its appreciated :)

OpenStudy (psymon):

Just to say, I was trying to prove convergence, not divergence. I used the p-series and saw that the original series was less than my p-series for all n. Maybe I'm missing something, but who knows o.o

OpenStudy (anonymous):

Who knows......and the battle continues :P

OpenStudy (psymon):

Fair enough. I remember when I was doing these I would forget whether the proof involved using another series or another limit. I would get the proofs mixed up and sometimes prove that a limit existed when I wanted to prove convergence instead. Oh well. You get through this part and you're good. This was honestly the most confusing partof calc II when i did it :P

OpenStudy (anonymous):

Im trying to make it through....only 3 more weekd sof this class

OpenStudy (psymon):

Yeah, you're at the end for sure. Do you need calc III, differential, or any of that stuff?

OpenStudy (anonymous):

i need both.

OpenStudy (psymon):

Gotcha. I'm taking differential and discrete math in a month. The website that linked me to openstudy in the first place is where I've been trying to preview D.Es. Might be worth a look. http://tutorial.math.lamar.edu/Classes/DE/DE.aspx

OpenStudy (psymon):

It has calc III on there as well, but I have to wait for the spring for that because of schedule conflict :/

OpenStudy (anonymous):

oh ok..ill check it out

OpenStudy (psymon):

: )